OpenStudy (jjuden):

A wooden plank is leaning against a wall. The bottom of the plank is 3 feet from the wall. Using the Sine ratio, find AC, the approximate length of the plank. Round to the nearest foot.

OpenStudy (jjuden):

heres my answer sin49= 3/AC AC=3/sin49 AC=3.96

OpenStudy (jjuden):